

********************************************************************
UARS WEEKLY STATUS REPORT - 20 Sep 2002
********************************************************************
***********
SOLSTICE
***********
During this week, SOLSTICE continued to operate in good health gathering
solar data on all available orbital tracking periods.
***********
SUSIM
***********
SUSIM has successfully performed its daily mid-resolution and MgII
solar scans during this past week for periods of predicted UARS
telemetry.
***********
ACRIM
***********
The ACRIM Science Team held discussions with FOT representatives and
obtained documentation pertaining to ACRIM2 power on/off sequences. A new
program of attempting to stimulate power-on operation is being designed
and will be implemented shortly.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 07 September 2002 (Orbit 60098, GMT Day 250)
through 13 September 2002 (Orbit 60203, GMT Day 256).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 251/1225 60121 POWER ON
PEM(Z/H) 252/1223 60136 POWER ON
HRDI 253/0342 60145 TO DAY ONLY MODE
HRDI 254/2156 60172 TO FULL TIME MODE
The beta angle ranged between the angles of 11.5 degrees to 34.6
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 57 degrees on 23 September 2002 (DOY 266).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.27 to 1.28 degrees C. Additional
battery performance data for each battery on 06 September and 13
September 2002 are provided in Appendix D.
The clock error ranged from -10.0 to 10.7 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
PEM = 01
SUSIM = 01
HALOE = 02
HRDI = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 00
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The UARS FOT is preparing for the integration of the ERBS FOT
beginning on 16 September 2002. ERBS hardware has begun making its
way over from Building 14 and will continue to flow in next week.
ERBS will take all of its day shift events next week from the
horseshoe in Building 32. We hope that the ERBS team has the same
success that we experienced in getting fully operational on the new
system.
The hard disk on the prototype will be repaired on 16 September 2002.
This will allow the FOT to test the patches that have been made by the
Eclipse development team. Once the patches have been tested a new
delivery of the software will be installed. This will most likely not
take place until the ERBS integration is complete.
MEETINGS
The FOT attended the UARS Monthly Status Meeting on 05 September 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 10 October 2002 (DOY 283).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 13 September 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
60105 250/1047 02-315 SUSIM POWER ON COMPLETE
60121 251/1225 02-317 HALOE POWER ON START
60126 251/2037 02-318 HALOE POWER ON COMPLETE
60135 252/1043 02-319 OBC/CLK CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
60136 252/1223 02-320 PEM POWER ON ZEPS/NEPS DAY-ONLY
60145 253/0342 02-321 HRDI COMMANDED TO DAY ONLY OPS
60172 254/2156 02-322 HRDI FROM DAY ONLY OPS TO FULLTIME
60188 255/2334 02-323 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 13 September 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 13 September 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
60135 252/1046 CLK ADJ FR -20.6 TO 0.0 MSEC/DAY
60188 255/2337 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 13 September 2002
06 September 2002 - GMT Day 249 Beta = 07.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -134.4/-156.8 -3.68/-4.94 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-100.8 +4.51/+3.87 +28.4/-12.8 25.4 1.020 19.5
3 +22.4/-22.4 +3.87/+2.92 +27.2/-12.8 25.4 1.010 19.0
13 September 2002 - GMT Day 256 Beta = 34.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -140.0/-162.4 -3.05/-4.31 +00.0/+00.0 08.3 0.000 0.0
2 +11.2/-123.2 +5.79/+4.51 +27.6/-15.2 24.5 1.040 20.9
3 +22.4/-67.2 +5.15/+3.87 +27.2/-15.2 24.3 1.040 20.6
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 13 September 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-045 FOT 59717 224/1458 14170 Generic Late Acquisition